# Yogesh Kulkarni  Pooyan Fazli

### v2

@@ 4.2 Temporal Advantage Shaping (TAS)
where $\lambda_{\text{TAS}}$ controls the amplitude of the weight variation. This function assigns the minimum weight of 1.0 to tokens at the midpoint of the sequence ($\tilde{t}=0.5$) and the maximum weight of $1.0+\lambda_{\text{TAS}}$ to tokens at the start and end ($\tilde{t}=0$ or 11). These weights define a token-specific shaped advantage: $A^{\text{TAS}}_{i,t}=w_{i,t}\cdot A_{i}$, where $i\in\{1,\dots,K\}$ indexes each of the $K$ candidate responses in the group. AVATAR’s loss is formed by replacing the standard GRPO advantage $A_{i}$ in Equation 2 with this shaped advantage. Reasoning sequence lengths, $L_{i}$, vary across completions, which can affect positional weighting. However, the TAS mechanism remains robust because the normalization in Equation 5 ensures that each token’s weight depends on its relative position $\tilde{t}$ within the sequence rather than its absolute position.
@@ 5.2 Ablation Studies
Table 5 validates the effectiveness of both the staged training curriculum and the TAS strategy. The curriculum shows consistent stage-wise improvements: Stage 11 RL strengthens visual reasoning over SFT, Stage 22 adds cross-modal alignment, and Stage 33 refines fine-grained localization, yielding progressive gains in reasoning ability. The advantage shaping ablation compares several weighting schemes: uniform (GRPO baseline), linear decay (planning-focused), linear incline (synthesis-focused), inverse parabolic (middle-focused), and our parabolic TAS (planning and synthesis-focused). Results confirm that extreme emphasis on either phase or the middle yields inferior results, while TAS achieves the best overall performance by jointly amplifying early planning and late synthesis steps.
@@ Appendix B Ablation Studies
Figure 7: Performance vs. on-policy/off-policy sample ratio across six benchmarks. AVATAR achieves optimal performance with 44-44 split (44 on-policy, 44 off-policy samples), demonstrating that balanced mixing prevents both policy drift from excessive off-policy data and sample inefficiency from pure on-policy training.
@@
Table 8: Ablation studies on our reward strategy across three training stages. Stage 11 uses $R_{\mathrm{acc}}$ and $R_{\mathrm{format}}$ for basic accuracy and format compliance. Stage 22 adds $R_{\mathrm{self}}$ (majority vote consensus) to the existing rewards, enabling consistency learning. Stage 33 incorporates $R_{\mathrm{judge}}$ (stepwise VLM evaluation) alongside all previous rewards, providing detailed reasoning feedback. VCRS maintains stable advantage baselines through moving averages, preventing vanishing advantages. Each cumulative addition demonstrates clear performance gains, with largest improvements on reasoning-heavy benchmarks (Video-Holmes: +4.5+4.5, MMVU: +5.4+5.4) where structured feedback proves most beneficial.
@@
To determine the optimal balance between on-policy and off-policy data, we analyze performance across different sample ratios within a fixed group size of eight (Figure 7). We observe a consistent trend across all benchmarks: performance peaks when using a balanced 44-44 split (44 on-policy, 44 off-policy samples). Using fewer off-policy samples (0-33) fails to fully mitigate the data inefficiency of a purely on-policy approach. On the other hand, over-reliance on off-policy samples (55-77) degrades performance due to instability, as the policy drifts too far from the older, “stale” data in the replay buffer. The 44-44 split therefore represents an empirical sweet spot for AVATAR, maximizing the data efficiency gains of our off-policy architecture while maintaining the training stability necessary for effective learning.
@@
Table 8 shows the cumulative impact of each reward component in AVATAR’s training pipeline. The baseline, which includes only accuracy and format rewards, offers limited learning signals, especially on complex reasoning tasks such as Video-Holmes and MMVU, where binary feedback proves insufficient. Introducing self-rewarding ($R_{\mathrm{self}}$) in Stage 22 enables consensus-based learning from the model’s own outputs, yielding moderate improvements on audio-visual tasks that benefit from internal consistency signals (e.g., +0.7+0.7 on OmniBench, +1.3+1.3 on MMVU). Adding the stepwise judge reward ($R_{\mathrm{judge}}$) in Stage 33 results in larger gains by assessing intermediate reasoning steps, particularly enhancing performance on fine-grained localization benchmarks (e.g., +0.8+0.8 on OmniBench, +1.3+1.3 on MMVU). Finally, incorporating VCRS further stabilizes training by maintaining non-zero advantage baselines via moving averages, allowing the full reward setup to achieve the highest observed performance (e.g., +0.8+0.8 on OmniBench, +0.9+0.9 on MMVU).
@@
In contrast, AVATAR shows a dip-and-recovery pattern that reflects the interplay of its key components. The initial drop in performance, especially on challenging benchmarks like OmniBench, Video-Holmes, and MMVU, occurs when the stratified replay buffer transitions from easy to hard samples, exposing the model to its most difficult failures. This performance dip would persist without TAS; however, the U-shaped temporal weighting in TAS provides focused learning signals that extract meaningful gradients from these difficult examples. By emphasizing the planning and synthesis phases that uniform credit assignment would dilute, TAS enables recovery and subsequent improvement. The model’s rebound and higher final performance confirm that our off-policy architecture transforms early failure into learning opportunities through repeated exposure, while TAS ensures each replayed experience yields maximum value. Notably, these gains emerge within just 10001000 iterations, underscoring AVATAR’s superior sample efficiency and training effectiveness.
@@ B.1 Prompt Templates
To support AVATAR’s training pipeline, we use two prompt templates for different components of our framework. The hint generation prompt (Figure 13) is used by our stratified replay buffer’s hint mechanism when the policy becomes stuck in local optima on challenging samples. This prompt instructs a teacher model to generate strategic hints without revealing the final answer, helping the policy escape stagnation while maintaining the learning challenge. The audio-visual localization judge prompt (Figure 14) helps our stepwise reasoning judge in Stage 33 training by providing granular evaluation criteria across four dimensions: grounding of audio cues, identification of visual objects, spatial localization accuracy, and caption correctness. This offers detailed feedback on cross-modal reasoning quality rather than binary success/failure signals.

### v3

@@ 4.2 Temporal Advantage Shaping (TAS)
where $\lambda_{\text{TAS}}$ controls the amplitude of the weight variation. This function assigns the minimum weight of 1.0 to tokens at the midpoint of the sequence ($\tilde{t}=0.5$) and the maximum weight of $1.0+\lambda_{\text{TAS}}$ to tokens at the start and end ($\tilde{t}=0$ or $1$). These weights define a token-specific shaped advantage: $A^{\text{TAS}}_{i,t}=w_{i,t}\cdot A_{i}$, where $i\in\{1,\dots,K\}$ indexes each of the $K$ candidate responses in the group. AVATAR’s loss is formed by replacing the standard GRPO advantage $A_{i}$ in Equation 2 with this shaped advantage. Reasoning sequence lengths, $L_{i}$, vary across completions, which can affect positional weighting. However, the TAS mechanism remains robust because the normalization in Equation 5 ensures that each token’s weight depends on its relative position $\tilde{t}$ within the sequence rather than its absolute position.
@@ 5.2 Ablation Studies
Table 5 validates the effectiveness of both the staged training curriculum and the TAS strategy. The curriculum shows consistent stage-wise improvements: Stage $1$ RL strengthens visual reasoning over SFT, Stage $2$ adds cross-modal alignment, and Stage $3$ refines fine-grained localization, yielding progressive gains in reasoning ability. The advantage shaping ablation compares several weighting schemes: uniform (GRPO baseline), linear decay (planning-focused), linear incline (synthesis-focused), inverse parabolic (middle-focused), and our parabolic TAS (planning and synthesis-focused). Results confirm that extreme emphasis on either phase or the middle yields inferior results, while TAS achieves the best overall performance by jointly amplifying early planning and late synthesis steps.
@@ Appendix B Ablation Studies
Figure 7: Performance vs. on-policy/off-policy sample ratio across six benchmarks. AVATAR achieves optimal performance with 44-44 split ($4$ on-policy, $4$ off-policy samples), demonstrating that balanced mixing prevents both policy drift from excessive off-policy data and sample inefficiency from pure on-policy training.
@@
Table 8: Ablation studies on our reward strategy across three training stages. Stage 11 uses $R_{\mathrm{acc}}$ and $R_{\mathrm{format}}$ for basic accuracy and format compliance. Stage $2$ adds $R_{\mathrm{self}}$ (majority vote consensus) to the existing rewards, enabling consistency learning. Stage 33 incorporates $R_{\mathrm{judge}}$ (stepwise VLM evaluation) alongside all previous rewards, providing detailed reasoning feedback. VCRS maintains stable advantage baselines through moving averages, preventing vanishing advantages. Each cumulative addition demonstrates clear performance gains, with largest improvements on reasoning-heavy benchmarks (Video-Holmes: +4.5+4.5, MMVU: +5.4+5.4) where structured feedback proves most beneficial.
@@
To determine the optimal balance between on-policy and off-policy data, we analyze performance across different sample ratios within a fixed group size of eight (Figure 7). We observe a consistent trend across all benchmarks: performance peaks when using a balanced 44-44 split ($4$ on-policy, $4$ off-policy samples). Using fewer off-policy samples (0-33) fails to fully mitigate the data inefficiency of a purely on-policy approach. On the other hand, over-reliance on off-policy samples (55-77) degrades performance due to instability, as the policy drifts too far from the older, “stale” data in the replay buffer. The 44-44 split therefore represents an empirical sweet spot for AVATAR, maximizing the data efficiency gains of our off-policy architecture while maintaining the training stability necessary for effective learning.
@@
Table 8 shows the cumulative impact of each reward component in AVATAR’s training pipeline. The baseline, which includes only accuracy and format rewards, offers limited learning signals, especially on complex reasoning tasks such as Video-Holmes and MMVU, where binary feedback proves insufficient. Introducing self-rewarding ($R_{\mathrm{self}}$) in Stage $2$ enables consensus-based learning from the model’s own outputs, yielding moderate improvements on audio-visual tasks that benefit from internal consistency signals (e.g., +0.7+0.7 on OmniBench, +1.3+1.3 on MMVU). Adding the stepwise judge reward ($R_{\mathrm{judge}}$) in Stage $3$ results in larger gains by assessing intermediate reasoning steps, particularly enhancing performance on fine-grained localization benchmarks (e.g., +0.8+0.8 on OmniBench, +1.3+1.3 on MMVU). Finally, incorporating VCRS further stabilizes training by maintaining non-zero advantage baselines via moving averages, allowing the full reward setup to achieve the highest observed performance (e.g., +0.8+0.8 on OmniBench, +0.9+0.9 on MMVU).
@@
In contrast, AVATAR shows a dip-and-recovery pattern that reflects the interplay of its key components. The initial drop in performance, especially on challenging benchmarks like OmniBench, Video-Holmes, and MMVU, occurs when the stratified replay buffer transitions from easy to hard samples, exposing the model to its most difficult failures. This performance dip would persist without TAS; however, the U-shaped temporal weighting in TAS provides focused learning signals that extract meaningful gradients from these difficult examples. By emphasizing the planning and synthesis phases that uniform credit assignment would dilute, TAS enables recovery and subsequent improvement. The model’s rebound and higher final performance confirm that our off-policy architecture transforms early failure into learning opportunities through repeated exposure, while TAS ensures each replayed experience yields maximum value. Notably, these gains emerge within just $1000$ iterations, underscoring AVATAR’s superior sample efficiency and training effectiveness.
@@ B.1 Prompt Templates
To support AVATAR’s training pipeline, we use two prompt templates for different components of our framework. The hint generation prompt (Figure 13) is used by our stratified replay buffer’s hint mechanism when the policy becomes stuck in local optima on challenging samples. This prompt instructs a teacher model to generate strategic hints without revealing the final answer, helping the policy escape stagnation while maintaining the learning challenge. The audio-visual localization judge prompt (Figure 14) helps our stepwise reasoning judge in Stage $3$ training by providing granular evaluation criteria across four dimensions: grounding of audio cues, identification of visual objects, spatial localization accuracy, and caption correctness. This offers detailed feedback on cross-modal reasoning quality rather than binary success/failure signals.